OpenStudy (anonymous):

use substitution method in system of equation to solve : 2x-y=10 2x-2y=4 a) (2.-1) b) ( 7,4) c) ( 2.-2) d) ( 8,6)

OpenStudy (mathmale):

Pretty please include ALL of the instructions when you post a problem. For example, in this case: "use substitution to solve the following system of linear equations." To get you started: Please solve the first equation for 2x.

OpenStudy (anonymous):

From the first equation, make y as the subject. That is 2x-y=10 2x-y + y=10 + y 2x = 10 + y 2x-10=10 + y - 10 2x - 10 = y That is y = 2x-10. Then substitute y = 2x-10 into the second equation to solve for x. Can you do it?

OpenStudy (anonymous):

i can try thanks

OpenStudy (anonymous):

Tell me what you have got when you finish :)

OpenStudy (anonymous):

sorry got distracted with something else . its (8,6)

OpenStudy (mathmale):

Hope it's not confusing for y ou, Geo, to have two different people suggesting that you try two different things. I suggested that you solve the first equation for 2x; RolyPoly suggested you solve it for y. Actually, either approach will work fine. Toss a coin and decide that way which to use on this particular problem. (RolyPoly's is the more traditional.)

OpenStudy (mathmale):

Substitute your x=8 and your y=6 back into the original equations. Are they both "true"? If so, your solution actually is a solution.

OpenStudy (anonymous):

its fine , it was right thanks both of y'all

OpenStudy (mathmale):

Good going! Congrats!

OpenStudy (anonymous):

thanks